OpenStudy (anonymous):

Write the equation y = -3x + 9 in standard form with integer coefficients

OpenStudy (anonymous):

i dont know how to do these can you teach me @mathmath333

OpenStudy (mathmath333):

only constant should be in RHS x and y variable should be in LHS like this \(\Large ax+by=c\)

OpenStudy (anonymous):

so what do i do

OpenStudy (mathmath333):

do u know LHS RHS ?

OpenStudy (anonymous):

um not really

OpenStudy (mathmath333):

LHS=left hand side RHS=right hand side any clue now

OpenStudy (anonymous):

oh okay so now that i now that what do i do with the equation

OpenStudy (mathmath333):

where is x variable in equation LHS or RHS

OpenStudy (anonymous):

rhs

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

ya right

OpenStudy (mathmath333):

so for bringing it , we need to shift it to LHS for that when we shift the term to other hand side only thr sign \(\huge \pm\) changes

OpenStudy (mathmath333):

example \(a=-b\\then\\ a+b=0\) or \(a=c\\then\\a-c=0\)

OpenStudy (anonymous):

When you bring something from RHS to LHS or LHS to RHS, then just reverse the sign as following: a) the positive term will become negative and vice-versa. b) the multiplication term will go onto divide and vice-versa.

OpenStudy (anonymous):

so we to get the x over to the lhs side right

OpenStudy (anonymous):

Like: \(x + 4 = y + 6\) Now suppose you want to bring \(+4\) to RHS, then bring it to RHS by changing its sign, thatn it + becomes _ now. \(x = y + 6 - 4\)

OpenStudy (anonymous):

See, in case of divide or multiplication, it is not that simple that you are thinking. For multiplication and divide, you should have only one term on RHS or LHS.

OpenStudy (anonymous):

Like \(x\) is in multiplication with \(-3\) there, you know to separate it from \(x\), the \(-3\) will go into division on LHS, but first take all the others to LHS, like you have \(+9\) there, so first move +9 and then divide.

OpenStudy (anonymous):

oh okay so how would i do the equation on top

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

you have : \(y = -3x + 9\)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\(-3x\) is a term, I mean as a whole it is a term only, so what is its sign?

OpenStudy (anonymous):

Like \(9\) as \(+\) sign there, \(y\) has also \(+\) sign, likewise \(3x\) has what sign?

OpenStudy (anonymous):

a negative

OpenStudy (anonymous):

yes, so negative on RHS will go by becoming positive on left? Right?

OpenStudy (anonymous):

See what is my post, I said, a negative term on RHS will go on LHS by changing its sign only, Plus becomes negative and negative becomes positive, getting?

OpenStudy (anonymous):

yes sorry ok so a number Neg on RHS will turn positive in LHS

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

\(y = \color{red}{-3x} + 9\) \(y \color{green}{+3x} = 9\), getting this?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

So anything more you want to do here?

OpenStudy (anonymous):

no thnxs sorry i took so long

OpenStudy (anonymous):

Now, standard equation to look nice, \(x\) term will be written first, and then \(y\) term, = and then constant. you must remember or note that: the coefficient of \(x\) must be positive.

OpenStudy (anonymous):

you see, 3x is positive as well as y is positive, so just rearrange that, write x term first and then y.

OpenStudy (anonymous):

\[\color{blue}{3x + y = 9}\]

OpenStudy (anonymous):

Look the coefficient of \(x\) is \(3\) which is \(+\), so that is what standard equation wants from us. :)

OpenStudy (anonymous):

so 3x+y=9

OpenStudy (anonymous):

Good, yes.